# Moving Least Squares without Quasi-Uniformity: A Stochastic Approach

### v1

@@ Abstract
Local Polynomial Regression (LPR) and Moving Least Squares (MLS) are closely related nonparametric estimation methods, developed independently in statistics and approximation theory. While statistical LPR analysis focuses on overcoming sampling noise under probabilistic assumptions, the deterministic MLS theory studies smoothness properties and convergence rates with respect to the fill-distance (a resolution parameter). Despite this similarity, the deterministic assumptions underlying MLS fail to hold under random sampling. We begin by quantifying the probabilistic behavior of the fill-distance $h_{n}$ and separation $\delta_{n}$ of an i.i.d. random sample. That is, for a distribution satisfying a mild regularity condition, $h_{n}\propto n^{-1/d}\log^{1/d}(n)$ and $\delta_{n}\propto n^{-1/d}$. We then prove that, for MLS of degree k1k\!-\!1, the approximation error associated with a differential operator $Q$ of order $|m|\leq k-1$ decays as $h_{n}^{\,k-|m|}$ up to logarithmic factors, establishing stochastic analogues of the classical MLS estimates. Additionally, We show that the MLS approximant is smooth with high probability. Finally, we apply the stochastic MLS theory to manifold estimation. Assuming that the sampled Manifold is $k$-times smooth, we show that the Hausdorff distance between the true manifold and its MLS reconstruction decays as $h_{n}^{k}$, extending the deterministic Manifold-MLS guarantees to random samples. This work provides the first unified stochastic analysis of MLS, demonstrating that—despite the failure of deterministic sampling assumptions—the classical convergence and smoothness properties persist under natural probabilistic models.
@@ 1.1 Background
In parallel, MLS was developed in computational mathematics as a meshfree method for function approximation Levin (1998). Like LPR, it employs weighted polynomial fitting, but its construction yields smooth shape functions with local support that enable continuous reconstruction across a domain. This property makes MLS especially suitable for geometric applications, such as surface reconstruction Levin (2004), solving partial differential equations Mirzaei (2015), and manifold approximation Sober and Levin (2020). The deterministic analyses of MLS rely on geometric sampling conditions, notably the notions of fill distance, the largest gap between sample points and the domain, and separation, the minimal distance between two distinct points. In this setting, the two quantities are comparable up to a constant factor, which provides a convenient framework for proving convergence and stability. Within this setting, Levin Levin (1998) demonstrated convergence rates for smooth functions, while Mirzaei Mirzaei (2015) extended the theory to derivatives. In particular, Mirzaei showed that if one applies a differential operator of order $m$ to both the MLS approximation and the true function, the difference between them decays like the fill distance raised to the power kmk-m, where $k$ is the smoothness order of the function. These results underscore the strength of MLS in preserving smoothness and geometric structure under deterministic sampling assumptions.
@@ 3 Main Results
$h_{n}=O_{P}(n^{-\frac{1}{d}}\ln^{\frac{1}{d}}(n))\quad\mbox{and}\quad\delta_{n}=\Omega_{P}(n^{-\frac{1}{d}}).$ (3)
@@ 4 Bridging the gap between deterministic and stochastic MLS
$\delta_{n}=\Omega_{P}(n^{-\frac{1}{d}}).$ (15)
@@
Recall that $P(B(x,\varepsilon))$ is the probability to have an $\varepsilon$-ball centered at $x$ with a sample in it. We’ve shown that this probability is bounded from above by $\bar{c}\cdot P_{U}(B(x,\varepsilon))$, and bounded from below by $\underline{c}\cdot P_{U}(B(x,\varepsilon))$. If we build the $\varepsilon$-balls around $x_{i}$, then the probability $P(\delta_{n}>\varepsilon)$ represents the probability that do not exist two samples in the same $\varepsilon$-ball. Then for n samples we get that
$P(\delta_{n}>\varepsilon)=\prod_{i=1}^{n-1}(1-i\cdot P(B(x_{i},\varepsilon)).$ (16)
@@
$P(\delta_{n}<\varepsilon)=1-\prod_{i=1}^{n-1}(1-i\cdot P(B(x_{i},\varepsilon)).$ (17)
@@
$1-\prod_{i=1}^{n-1}(1-i\cdot P\left(B(x_{i},\varepsilon\right)))\leq 1-\prod_{i=1}^{n-1}(1-i\cdot\bar{c}\cdot P_{U}(B(x_{i},\varepsilon))),$ (18)
@@
Similarly to Lemma 7 we now express $\varepsilon$ in terms of n, and for any $\rho_{\delta}\in(0,1)$ if we let
$1-\left(1-nc^{\prime}\varepsilon^{d}\right)^{n}\leq\rho_{\delta},$ (23)
then by simple algebra we obtain that exist $K_{\delta}$, a real positive constant, such that
$\varepsilon\leq\left(\frac{1-(1-\rho_{\delta})^{1/n}}{c^{\prime}n}\right)^{1/d}=K_{\delta}\cdot n^{-\frac{1}{d}}.$ (24)
Now we can rewrite Equation (22) as
$P(\delta_{n}<\varepsilon)\leq\rho_{\delta}$
which is equivalent to
$P\left(\left|\frac{\delta_{n}}{\varepsilon}\right|<1\right)\leq\rho_{\delta}.$
If we substitute the upper bound of $\varepsilon$ from Equation (24), the inequality remains valid. Hence, for any $\rho_{\delta}>0$, exists a finite $N_{\delta}>0$ such that
$P\left(\left|\frac{\delta_{n}}{K_{\delta}\cdot n^{-\frac{1}{d}}}\right|<1\right)\leq\rho_{\delta},\,\,\forall n>N_{\delta}$
or, equivalently, $\delta_{n}=\Omega_{P}(n^{-\frac{1}{d}})$. ∎
@@
Since $h_{n}=O_{P}(n^{-\frac{1}{d}}\ln^{\frac{1}{d}}(n))$ and $\delta_{n}=\Theta_{P}(n^{-\frac{1}{d}})$ we get that the quasi-uniform property does not hold (see Definition 2).
@@ 5 Proof of Theorem 2
We want to prove that $\mathscr{A}_{n}(x)$ is uniformly positive definite, i.e., its smallest eigenvalue is bounded away from zero independently of $n$, given 1, 43 and under Lemma 3.

### v2

@@ Abstract
Local Polynomial Regression (LPR) and Moving Least Squares (MLS) are closely related nonparametric estimation methods, developed independently in statistics and approximation theory. While statistical LPR analysis focuses on overcoming sampling noise under probabilistic assumptions, the deterministic MLS theory studies smoothness properties and convergence rates with respect to the fill-distance (a resolution parameter). Despite this similarity, the deterministic assumptions underlying MLS fail to hold under random sampling. We begin by quantifying the probabilistic behavior of the fill-distance $h_{n}$ and separation $\delta_{n}$ of an i.i.d. random sample. That is, for a distribution satisfying a mild regularity condition, $h_{n}\propto n^{-1/d}\log^{1/d}(n)$ and $\delta_{n}\propto n^{-2/d}$. We then prove that, for MLS of degree k1k\!-\!1, the approximation error associated with a differential operator $Q$ of order $|m|\leq k-1$ decays as $h_{n}^{\,k-|m|}$ up to logarithmic factors, establishing stochastic analogues of the classical MLS estimates. Additionally, We show that the MLS approximant is smooth with high probability. Finally, we apply the stochastic MLS theory to manifold estimation. Assuming that the sampled Manifold is $k$-times smooth, we show that the Hausdorff distance between the true manifold and its MLS reconstruction decays as $h_{n}^{k}$, extending the deterministic Manifold-MLS guarantees to random samples. This work provides the first unified stochastic analysis of MLS, demonstrating that—despite the failure of deterministic sampling assumptions—the classical convergence and smoothness properties persist under natural probabilistic models.
@@ 1.1 Background
In parallel, MLS was developed in computational mathematics as a meshfree method for function approximation Levin (1998). Like LPR, it employs weighted polynomial fitting, but its construction yields smooth shape functions with local support that enable continuous reconstruction across a domain. This property makes MLS especially suitable for geometric applications, such as surface reconstruction Levin (2004), solving partial differential equations Mirzaei (2015), and manifold approximation Sober and Levin (2020). The deterministic analyses of MLS rely on geometric sampling conditions, notably the notions of fill distance (sometimes referred to as the covering radius), the largest gap between sample points and the domain, and separation, the minimal distance between two distinct points. In this setting, the two quantities are comparable up to a constant factor, which provides a convenient framework for proving convergence and stability. Within this setting, Levin Levin (1998) demonstrated convergence rates for smooth functions, while Mirzaei Mirzaei (2015) extended the theory to derivatives. In particular, Mirzaei showed that if one applies a differential operator of order $m$ to both the MLS approximation and the true function, the difference between them decays like the fill distance raised to the power kmk-m, where $k$ is the smoothness order of the function. These results underscore the strength of MLS in preserving smoothness and geometric structure under deterministic sampling assumptions. More recently, Krieg and Sonnleitner Krieg and Sonnleitner (2024) relaxed the reliance on fill distance by showing that random points on convex domains are asymptotically optimal for Sobolev approximations. This suggests that the large gaps characteristic of stochastic sampling do not prevent optimal average convergence rates.
@@ 3 Main Results
$h_{n}=O_{P}(n^{-\frac{1}{d}}\ln^{\frac{1}{d}}(n))\quad\mbox{and}\quad\delta_{n}=\Omega_{P}(n^{-\frac{2}{d}}).$ (3)
@@ 4 Bridging the gap between deterministic and stochastic MLS
Remark 3.
A similar upper bound for $h_{n}$ was first established in Reznikov and Saff (2016). We include this proof for completeness.
@@
$\delta_{n}=\Omega_{P}(n^{-\frac{2}{d}}).$ (15)
Remark 4.
A version of the bound in Lemma 15 was previously established by Brauchart et al. (2018) for the specific case of a sphere under a uniform distribution. Our result generalizes this bound to any compact domain in $\mathbb{R}^{n}$ and relaxes the distribution requirement to almost uniform distributions.
@@
Recall that $P(B(x,\varepsilon))$ is the probability to have an $\varepsilon$-ball centered at $x$ with a sample in it. We’ve shown that this probability is bounded from above by $\bar{c}\cdot P_{U}(B(x,\varepsilon))$, and bounded from below by $\underline{c}\cdot P_{U}(B(x,\varepsilon))$. Note that $P(\delta_{n}>\varepsilon)$ represents the probability that no two samples from $X$ occupy the same $\varepsilon$-ball around either of them. Thus, considering $\varepsilon$-balls around all samples and using the fact that these balls may not be disjoint, we get the following bound
$P(\delta_{n}>\varepsilon)\geq\prod_{i=1}^{n-1}(1-\sum_{j=1}^{i}P(B(x_{j},\varepsilon)).$ (16)
@@
$P(\delta_{n}<\varepsilon)\leq 1-\prod_{i=1}^{n-1}(1-\sum_{j=1}^{i}P(B(x_{j},\varepsilon)).$ (17)
@@
$1-\prod_{i=1}^{n-1}(1-\sum_{j=1}^{i}P(B(x_{j},\varepsilon))\leq 1-\prod_{i=1}^{n-1}(1-i\cdot\bar{c}\cdot P_{U}(B(x_{i},\varepsilon))),$ (18)
@@
Fix any $\rho_{\delta}\in(0,1)$. We seek to find a sequence $\varepsilon(n)$ such that the right-hand side of (22) is bounded by $\rho_{\delta}$. Let $\varepsilon(n)$ be defined by setting:
$1-(1-nc^{\prime}\varepsilon^{d})^{n}=\rho_{\delta}.$ (23)
Solving for $\varepsilon(n)$ yields:
$\varepsilon(n)=\left(\frac{1-(1-\rho_{\delta})^{1/n}}{nc^{\prime}}\right)^{1/d}.$ (24)
To obtain a clean bound for (24), we observe that $(1-\rho_{\delta})^{1/n}=\exp\left(\frac{1}{n}\log(1-\rho_{\delta})\right)$. Applying the inequality $e^{-x}\geq 1-x$ for $x\geq 0$, where $x=\frac{1}{n}\log\left(\frac{1}{1-\rho_{\delta}}\right)$, we have:
$(1-\rho_{\delta})^{1/n}\geq 1-\frac{\log(1/(1-\rho_{\delta}))}{n}.$ (25)
Rearranging this inequality gives:
$1-(1-\rho_{\delta})^{1/n}\leq\frac{\log(1/(1-\rho_{\delta}))}{n}.$ (26)
Substituting this into (24), we define the constant $K_{\delta}$ as:
$K_{\delta}:=\left(\frac{\log(1/(1-\rho_{\delta}))}{c^{\prime}}\right)^{1/d}.$ (27)
It follows that for all $n\geq 1$:
$\varepsilon(n)\leq K_{\delta}\,n^{-2/d}.$ (28)
Consequently, substituting this bound back into (22), we obtain:
$P\left(\delta_{n}<K_{\delta}\,n^{-2/d}\right)\leq\rho_{\delta},\quad\forall n\geq 1.$ (29)
By the definition of order in probability for lower bounds, this strictly implies:
$\delta_{n}=\Omega_{P}(n^{-2/d}).$ (30)
@@
Since $h_{n}=O_{P}(n^{-\frac{1}{d}}\ln^{\frac{1}{d}}(n))$ and $\delta_{n}=\Omega_{P}(n^{-\frac{2}{d}})$ we get that the quasi-uniform property does not hold (see Definition 2).
@@ 5 Proof of Theorem 2
We want to prove that $\mathscr{A}_{n}(x)$ is uniformly positive definite, i.e., its smallest eigenvalue is bounded away from zero independently of $n$, given 1, 49 and under Lemma 3.